# Post-detection inference for sequential changepoint localization

### v3

@@ 6 Experiments
Comparison with [36]: We now compare our method with Ch. 3.5 of [36] on the Gaussian mean-change, having the (unknown) changepoint at $T=100$ or 500500500. We recall that, unlike our general nonasymptotic guarantees for any $\mathcal{A}$, Wu’s method only has asymptotic approximate guarantees in specific parametric settings, and only for CUSUM procedure [18], $\tau^{\prime}_{\text{cusum}}:=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}T_{j}\geq d\right\},$ where $T_{n}=\max\{0,T_{n-1}+X_{n}\}$ and $d$ is the threshold. We used $d=8.59$ and 7.567.567.56 for ${N}(-0.25,1)$ vs. ${N}(0.25,1)$ and ${N}(-0.3,1)$ vs. ${N}(0.3,1)$ respectively. We implement Wu’s confidence set of the form $V_{c}\cup[L_{s},\hat{\nu})$, where $\hat{\nu}=\max\{k<\tau:T_{k}=0\}$, $V_{c}=\{k\geq\hat{\nu}:T_{k}\leq c\}$, and $L_{k}$ denote the $k$-th last zero point of $T_{n}$ counted backward starting from $L_{0}:=\hat{\nu}$, with their proposed approximations (for the case when $\theta_{1}=-\theta_{0}$) $s=\log(\alpha)\times(1/(\sqrt{2}\theta_{1})+0.088)$ and $c=-\log(1-\sqrt{1-\alpha})/(2\theta_{1})-0.583$ with $\alpha=0.05$. Table 6.11 reports the average size and coverage rates (for target conditional coverage 0.950.950.95) across 500500500 runs.
@@
In Settings I, II, and III, both of our proposed methods — universal (Section 3) and adaptive (Section 4) — are applicable. For nonparametric settings (IV and V), the only valid method currently known is the universal method. Prior works are valid (asymptotically) only for Setting I. As shown in Table 6.2 for known pre- and post-change distributions (Setting I), the adaptive method consistently outperforms the universal method, which tends to be conservative. Also, our adaptive method is the only method we know that remains valid under dependence (see experimental results in Supplement G.1). However, for the composite settings (II and III), the universal method outperforms the adaptive method. Moreover, the universal method is also significantly more efficient — it runs in $O(T)$ time (assuming $M_{t}$ can be computed in $O(1)$ time), while the adaptive method incurs a higher computational cost of $O(T^{2}B)$ (assuming $\mathcal{A}$ runs in $O(T)$ time and $M_{t}$ can be computed in $O(1)$ time). See Supplement Sections D, E, and G for further experiments, involving confidence intervals for pre- and post-change parameters, detection procedures with bounded PFA, Poisson rate change problems, settings with dependent data, etc.
@@ 7 Conclusion
Omitted proofs can be found in Appendix A. We provide the implementation details of Algorithms 2 and 3 in Appendices B and C respectively, along with two concrete examples in both sections. Experimental results for the confidence intervals for pre- and post-change parameters are in Appendix D. Experimental results for a different implementation method are in Appendix E. Simpler algorithms for PFA-controlling change detectors are presented in Appendix F. Some additional experiments with dependent (Markov) pre- and post-change data, with detection procedures with bounded PFA, and with Poisson rate change problems are in Appendix G. Some heuristic algorithms are presented in Appendices H and I, pointing at open avenues for further theoretical work.
@@ Appendix A Omitted proofs
First note the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T),$ since the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
Let $\mathcal{C}\subseteq\{1,\dots,\tau\}$ be a confidence set satisfying $\mathbb{P}_{F_{0},t,F_{1}}(t\in\mathcal{C}\mid\tau\geq t)\geq 1-\alpha$, for all $t\in\mathbb{N}$. Since $\tau$ be a stopping time, $[\tau<t]=\displaystyle\cup_{i=1}^{t-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq t\}$. This implies $\mathbb{P}_{F_{0},\infty}(\tau\geq t)=\mathbb{P}_{F_{0},\infty}(\tau\geq t).$ Therefore,
@@
$\leq\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t),\leavevmode\nobreak\ \forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$
@@
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\}\mid\tau\geq T)=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C},\tau\geq T)}{\mathbb{P}_{F_{0},T,F_{1}}(\tau\geq T)}=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\geq 1-\alpha.$
@@
$\mathbb{P}\left(\exists n\in\mathbb{N}:R_{n}\geq 1/\alpha\right)\leq\alpha,\forall\mathbb{P}\in\mathcal{P}.$ (A.1)
@@
$=P_{F_{1}}\left(\exists n\geq t,R^{(t)}_{n}\geq 2/\alpha\right)\leq\alpha/2.$
@@
From Theorem 3.1, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]\leq\alpha r_{t}$, since $r_{t}$ and $M_{t}$ are independent of each other. Hence,
$\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}]=\mathbb{E}[\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]]\leq\mathbb{E}[\alpha r_{t}]=\alpha\mathbb{P}_{F_{0},\infty}[\tau\geq t],\forall t\in\mathbb{N}.$ Now,
@@
where we twice used the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T)$, which holds because $\tau$ is a stopping time and so the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$. So, we obtain
@@
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\mathbb{P}_{F_{0},T,F_{1}}(M_{T}>2/(\alpha r_{T}))}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\geq 1-\alpha.$ (A.2)
@@
We only prove the case when $\theta_{0}^{*}=\sup\Theta_{0}\leq\inf\Theta_{1}$ and $\tau$ is the CUSUM type stopping time $\tau_{\text{wcs-ripr}}$ defined in (C.12). Proofs for the other cases, when $\theta_{0}^{*}=\inf\Theta_{0}\geq\sup\Theta_{1}$ and/or $\tau$ is the SR-type stopping time (C.13), follow analogously.
$\mathbb{P}_{\theta,\infty}[\tau_{\text{wcs-ripr}}\geq t]$ $=\mathbb{P}_{\theta,\infty}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta_{1}\in\Theta_{1}}\prod_{i=j}^{k}\frac{f_{\theta_{1}}(X_{i})}{f_{\theta_{0}^{*}}(X_{i})}dW(\theta_{1})<A\right].$
Since the family $\{F_{\theta}\}$ satisfies the MLR property in the statistic $T(x)$, and since $\frac{f_{\theta_{1}}(x)}{f_{\theta_{0}^{*}}(x)}$ is non-decreasing in $T(x)$, it follows that the random variables $\frac{f_{\theta_{1}}(X)}{f_{\theta_{0}^{*}}(X)}$ are stochastically increasing in $\theta$. That is, for $X\sim F_{\theta}$ and $X^{\prime}\sim F_{\theta^{\prime}}$ with $\theta\leq\theta^{\prime}\leq\sup\Theta_{0}$, we have $T(X)\preceq_{\text{st}}T(X^{\prime})$ (see e.g., [16, Chapter 3.4]) and hence,
$\frac{f_{\theta_{1}}(X)}{f_{\theta_{0}^{*}}(X)}\preceq_{\text{st}}\frac{f_{\theta_{1}}(X^{\prime})}{f_{\theta_{0}^{*}}(X^{\prime})}.$
By Theorem 1.A.2 of [24], the stochastic ordering implies that there exist a random variable $Z_{\theta_{1}}$ and functions $\psi_{\theta_{1},1}$ and $\psi_{\theta_{1},2}$ such that $\psi_{\theta_{1},1}(z)\leq\psi_{\theta_{1},2}(z)$ for all $z$ and $\frac{f_{\theta_{1}}(X)}{f_{\theta_{0}^{*}}(X)}\stackrel{{\scriptstyle d}}{{=}}\psi_{\theta_{1},1}(Z_{\theta_{1}})$ and $\frac{f_{\theta_{1}}(X^{\prime})}{f_{\theta_{0}^{*}}(X^{\prime})}\stackrel{{\scriptstyle d}}{{=}}\psi_{\theta_{1},2}(Z_{\theta_{1}})$. Therefore,
$\mathbb{P}_{\theta,\infty}[\tau_{\text{wcs-ripr}}\geq t]$ $=\mathbb{P}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta_{1}\in\Theta_{1}}\prod_{i=j}^{k}\psi_{\theta_{1},1}(Z_{\theta_{1},i})dW(\theta_{1})<A\right]$
$\geq\mathbb{P}\left[\displaystyle\max_{1\leq k\leq t}\max_{1\leq j\leq k}\int_{\theta_{1}\in\Theta_{1}}\prod_{i=j}^{k}\psi_{\theta_{1},2}(Z_{\theta_{1},i})dW(\theta_{1})<A\right]$
$=\mathbb{P}_{\theta^{\prime},\infty}[\tau_{\text{wcs-ripr}}\geq t].$

### v4

@@ 6 Experiments
Huber-robust (nonparametric) pre- and post-change (Setting VI):
We perform experiments for the robust Gaussian mean-change problem under the $\epsilon$ contamination model, with the true changepoint at $T=100,500$. The pre- and post-change classes are $\epsilon$ neighbourhoods around $N(\mu_{0},1)$ and $N(\mu_{1},1)$, respectively: $\mathcal{P}_{i}=\{(1-\epsilon)N(\mu_{i},1)+\epsilon F:F\in\mathcal{M}\};\penalty 10000\ \penalty 10000\ i=0,1.$ In our experiment, $\mu_{0}=0,\mu_{1}=1$ and $\epsilon=0.01,0.001$. The true (unknown) pre- and post-change distributions are $(1-\epsilon)\times N(0,1)+\epsilon\times\text{Cauchy}(-1,10)$ and $(1-\epsilon)\times N(1,1)+\epsilon\times\text{Cauchy}(-1,10)$ respectively. We use Huber’s LFD (which are also log-optimal e-values [24]) to construct a CUSUM-style change-detector as below:
$\tau=\inf\left\{t\in\mathbb{N}:\max_{1\leq j\leq t}\prod_{i=j}^{t}\pi(X_{i})\geq A\right\},$ (6.7)
where $\pi(x)=\min\{\max\{\frac{f_{1}(x)}{f_{0}(x)},c^{\prime}\},c^{\prime\prime}\}$ for constants $c^{\prime\prime}>c^{\prime}\geq 0$ as mentioned in [13, 24]. We set the threshold at $A=1000$. $\hat{T}$ is as defined in (3.3) with the same e-process, $\prod_{i=1}^{t}\pi(X_{i})$, as mentioned above. After detection, we construct our universal confidence sets (3.9) for $T$, with $R_{n}^{(t)}=1/\prod_{i=t}^{n}\pi(X_{i})$, for $n\geq t$ $S_{n}^{(t)}=\prod_{i=n}^{t-1}\pi(X_{i})$, for $n<t$ and $\alpha=0.05$. We report the average size and coverage rates across $200$ independent runs in Table 6.9. We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection).
Table 6.9: Setting VI: Huber-robust (nonparametric) pre- and post-change
T $\epsilon$
Conditional
coverage
Marginal
coverage
Size
$|\hat{T}-T|$
Delay
100 0.01 0.98 0.98 16.74 3.23 18.38
100 0.001 0.96 0.96 15.57 3.18 13.96
500 0.01 0.97 0.93 16.18 3.87 16.58
500 0.001 0.97 0.92 15.13 3.25 14.44
[…]
@@
Sentiment change in language data: We consider the Stanford Sentiment Treebank (SST-2) dataset of movie reviews labeled with a binary sentiment [31]. We want to localize the change from $60\%$ positive reviews to $40\%$ positive reviews. In our experimental setup, the true changepoint is at 500; the pre-change data are i.i.d. draws of $75\%$ positive reviews, and the post-change data are i.i.d. draws of $25\%$ negative reviews (these are all unknown to our algorithm). We construct a change detection method and a point estimate $\hat{T}$, using a pre-trained DistilBERT-base model fine-tuned on the uncased SST-2 [25], which is a binary sentiment classifier that maps each review to either 0 (negative) or $1$ (positive), thereby producing a sequence of Bernoulli observations. After obtaining the binary sequence and $\hat{T}$, we construct the test statistic (3.4) using the likelihood ratio process for $\text{Bern}(0.4)$ against $\text{Bern}(0.6)$ (forward) and for $\text{Bern}(0.6)$ against $\text{Bern}(0.4)$ (backward). Note that given the change detector (and as a byproduct, the sequence of Bernoulli) and $\hat{T}$, it is simple to compute (3.4) in $O(T)$ time per $t$. So, the confidence set (3.9) takes total $O(T^{2})$ time. Fig. 6.11 provides a visualization.
Figure 6.11: Vertical red dashed lines show the point estimate. The confidence set (3.9) is marked in red points. The true changepoint $T=500$ is marked with the black line. Texts are plotted from time $t=470$ upto detection time, $\tau=522$.
@@
In Settings I, II, and III, both of our proposed methods — universal (Section 3) and adaptive (Section 4) — are applicable. For nonparametric settings (IV, V, and VI), the only valid method currently known is the universal method. Prior works are valid (asymptotically) only for Setting I. As shown in Table 6.2 for known pre- and post-change distributions (Setting I), the adaptive method consistently outperforms the universal method, which tends to be conservative. Also, our adaptive method is the only method we know that remains valid under dependence (see experimental results in Supplement H.1). However, for the composite settings (II and III), the universal method outperforms the adaptive method. Moreover, the universal method is also significantly more efficient — it runs in $O(T^{2})$ time (assuming $M_{t}$ can be computed in $O(T)$ time), while the adaptive method incurs a higher computational cost of $O(T^{3}B)$ (assuming $\mathcal{A}$ runs in $O(T)$ time and $M_{t}$ can be computed in $O(T)$ time). See Supplement Sections E, F, and H for further experiments, involving confidence intervals for pre- and post-change parameters, detection procedures with bounded PFA, settings with dependent data, etc.
@@ 7 Conclusion
Omitted proofs can be found in Appendix A. We provide the implementation details of Algorithm 1 in Appendix B and of Algorithms 2 and 3 in Appendices C and D respectively, along with two concrete examples in both sections. Experimental results for the confidence intervals for pre- and post-change parameters are in Appendix E. Experimental results for a different implementation method are in Appendix F. Simpler algorithms for PFA-controlling change detectors are presented in Appendix G. Some additional experiments with dependent (Markov) pre- and post-change data, with detection procedures with bounded PFA, and with Poisson rate change problems are in Appendix H. Some heuristic algorithms are presented in Appendices I and J, pointing at open avenues for further theoretical work.
@@ Appendix A Omitted proofs
First note the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T),$ since the event $\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
Let $\mathcal{C}\subseteq\{1,\dots,\tau\}$ be a confidence set satisfying $\mathbb{P}_{F_{0},t,F_{1}}(t\in\mathcal{C}\mid\tau\geq t)\geq 1-\alpha$, for all $t\in\mathbb{N}$. Since $\tau$ be a stopping time, $\tau<t]=\displaystyle\cup_{i=1}^{t-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq t\}$. This implies $\mathbb{P}_{F_{0},\infty}(\tau\geq t)=\mathbb{P}_{F_{0},\infty}(\tau\geq t).$ Therefore,
@@
$\leq\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t),\penalty 10000\ \forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$
@@
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C},\tau\geq T)}{\mathbb{P}_{F_{0},T,F_{1}}(\tau\geq T)}=\frac{\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\geq 1-\alpha.$
@@
$\mathbb{P}\left(\exists n\in\mathbb{N}:R_{n}\geq 1/\alpha\right)\leq\alpha,\forall\mathbb{P}\in\mathcal{P}.$ (S1)
@@
$=\mathbb{P}_{F_{0},t,F_{1}}\left(\exists n\geq t,R^{(t)}_{n}\geq 2/\alpha\right)\leq\alpha/2.$
@@
Observe that
@@
where we twice used the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T)$, which holds because $\tau$ is a stopping time and so the event $\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
From Theorem 3.1, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]\leq\alpha r_{t}$, since $r_{t}$ and $M_{t}$ are independent of each other. Now, since $r_{t}$ is an unbiased estimator of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}]=\mathbb{E}[\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]]\leq\mathbb{E}[\alpha r_{t}]=\alpha\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, $\forall t\in\mathbb{N}.$ So, we obtain
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1$- $\frac{\mathbb{P}_{F_{0},T,F_{1}}(M_{T}>2/(\alpha r_{T}))}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\geq 1-\alpha.$
Proof of Theorem 3.3.
We first decompose the length of the confidence set into contributions from the pre- and post-change regions, and then control each contribution separately.
Decomposition of the confidence set.
Since $M_{t}=\frac{L_{\hat{T}}^{\tau}}{L_{t}^{\tau}}$ and $\hat{T}=\arg\max_{1\leq i\leq\tau}L_{i}^{\tau}$ and hence, whenever $\tau\geq T>t,$
$M_{t}\geq\frac{L_{{T}}^{\tau}}{L_{t}^{\tau}}=\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}.$
Similarly, whenever $T<t\leq\tau,$ $M_{t}=\frac{L_{\hat{T}}^{\tau}}{L_{t}^{\tau}}\geq\frac{L_{{T}}^{\tau}}{L_{t}^{\tau}}=\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}$. Therefore,
$|{\mathcal{C}}|\times\mathds{1}(\tau\geq T)$
$\leq\sum_{t=1}^{{T-1}}\mathds{1}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)+1+\sum_{t=T+1}^{{\tau}}\mathds{1}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\leq 2/(\alpha r_{t})\right)$
$=\sum_{t=1}^{{T-1}}\mathds{1}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)+1+\sum_{t=T+1}^{{\infty}}\mathds{1}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\leq 2/(\alpha r_{t}),t\leq\tau\right).$
We also have
$|{\mathcal{C}}|\leq\sum_{t=1}^{{T-1}}\mathds{1}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)+1+|\tau-T|.$ (S2)
So, taking the conditional expectation on both sides and combining both bounds, we get
$\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq 1+\sum_{t=1}^{{T-1}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\mid\tau\geq T\right)$
@@
$+\min\left\{\sum_{t=T+1}^{{\infty}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\leq 2/(\alpha r_{t}),t\leq\tau\mid\tau\geq T\right),\Delta_{F_{0},T,F_{1}}\right\}.$ (S3)
Controlling the Pre-Change Tail.
Fix some $s\in(0,1)$. Then, using Markov’s equality,
$\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)$ $\leq\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{T})\right)\quad\text{[Since }r_{t}\leq r_{T}]$
$=\mathbb{E}\left(\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\left(\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\right)^{s}\geq(\alpha r_{T})^{s}/2^{s}\mid r_{T}\right)\right)$
$\leq\mathbb{E}\left(\left(\frac{2}{\alpha r_{T}}\right)^{s}\mathbb{E}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\left(\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\right)^{s}\right)\right)$
$=\mathbb{E}\left(\left(\frac{2}{\alpha r_{T}}\right)^{s}\left(\mathbb{E}_{F_{0}}\left(\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\right)^{s}\right)^{T-t}\right)$
$=\mathbb{E}\left(\frac{2}{\alpha r_{T}}\right)^{s}\rho_{0}(s)^{T-t}.$
This holds for all $s\in(0,1)$ and hence for $s=s_{0}.$ As noted earlier, $\rho_{0}(s_{0})<1$. So, we have
$\sum_{t=1}^{{T-1}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\mid\tau\geq T\right)$ $\leq\frac{1}{p_{T}}\sum_{t=1}^{{T-1}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)$
$\leq\mathbb{E}\left(\frac{2}{\alpha r_{T}}\right)^{s}\frac{1}{p_{T}}\sum_{t=1}^{{T-1}}\rho_{0}(s)^{T-t}$
$\to\frac{2^{s}}{\alpha^{s}p_{{T}}^{s+1}}\times\frac{\rho_{0}(s)-\rho_{0}(s)^{T-1}}{1-\rho_{0}(s)},$
as $N\to\infty.$ To justify the convergence, we note that $\left(\frac{2}{\alpha r_{T}}\right)^{s}$ is uniformly integrable as a sequence in $N\in\mathbb{N}$ (follows from Lemma A.2), and by the strong law of large numbers, $\left(\frac{2}{\alpha r_{T}}\right)^{s}\to\left(\frac{2}{\alpha p_{T}}\right)^{s}$ almost surely. Since the above holds for any $s\in(0,1)$, it also holds for $s=s_{0}.$
Deriving the first bound.
So, taking the conditional expectation on both sides of (S2), we get
$\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq 1+\sum_{t=1}^{{T-1}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\mid\tau\geq T\right)+\Delta_{F_{0},T,F_{1}}.$
Therefore, it follows from the above calculation that
$\limsup_{N\to\infty}\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq\frac{2^{s_{0}}}{\alpha^{s_{0}}p_{{T}}^{s_{0}+1}}\times\frac{\rho_{0}(s_{0})-\rho_{0}(s_{0})^{T-1}}{1-\rho_{0}(s_{0})}+1+\Delta_{F_{0},T,F_{1}}.$ (S4)
Controlling the “Post-Change” Tail when $\tau$ is sensitive.
Similarly, for $t\geq T+1$, we have
$\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\leq 2/(\alpha r_{t}),t\leq\tau\right)$
$\leq\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\leq 2/(\alpha r_{\tau})\right)$
$\leq\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\leq 2/(\alpha r_{\tau}),r_{\tau}>\rho_{1}(s)^{t-T}\right)+\mathbb{P}_{F_{0},T,F_{1}}[r_{\tau}\leq\rho_{1}(s)^{t-T}]$
$\leq\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\left(\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\right)^{s}\geq(\alpha\rho_{1}(s)^{t-T})^{s}/2^{s}\right)+\mathbb{P}_{F_{0},T,F_{1}}[r_{\tau}\leq\rho_{1}(s)^{t-T}]$
$\leq\left(\frac{2}{\alpha\rho_{1}(s)^{t-T}}\right)^{s}\mathbb{E}_{F_{1}}\left(\prod_{i=T}^{{t}-1}\left(\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\right)^{s}\right)+\mathbb{P}_{F_{0},T,F_{1}}[r_{\tau}\leq\rho_{1}(s)^{t-T}]$
$\leq\left(\frac{2}{\alpha\rho_{1}(s)^{t-T}}\right)^{s}\rho_{1}(s)^{t-T}+\rho_{1}(s)^{t-T}.$
The last step follows from the assumption that $\tau$ is sensible (recall that if $\tau$ is sensitive, then $\tau_{j}$ is stochastically larger than $\tau$) and from Lemma A.3. Also, as we noted before, $\rho_{1}(s_{1})<1.$ Therefore,
$\sum_{t=T+1}^{{\infty}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\leq 2/(\alpha r_{t}),\tau\geq t\mid\tau\geq T\right)$
$\leq\frac{1}{p_{T}}\sum_{t=T+1}^{{\infty}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\leq 2/(\alpha r_{t}),\tau\geq t\right)$
$\leq\frac{1}{p_{T}}\sum_{t=T+1}^{{\infty}}\left(\left(\frac{2}{\alpha}\right)^{s}\rho_{1}(s)^{s(t-T)}+\rho_{1}(s)^{t-T}\right)$
$\leq\frac{1}{p_{T}}\left(\left(\frac{2}{\alpha}\right)^{s}\times\frac{\rho_{1}(s)^{s}}{1-\rho_{1}(s)^{s}}+\frac{\rho_{1}(s)}{1-\rho_{1}(s)}\right).$
Since the above holds for any $s\in(0,1)$, it also holds for $s=s_{1}.$
Deriving the final bound.
Finally, it follows from (A) and the above calculations that
$\limsup_{N\to\infty}\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq\frac{2^{s_{0}}}{\alpha^{s_{0}}p_{{T}}^{s_{0}+1}}\times\frac{\rho_{0}(s_{0})-\rho_{0}(s_{0})^{T-1}}{1-\rho_{0}(s_{0})}+1$
$+\min\left\{\frac{1}{p_{T}}\left(\left(\frac{2}{\alpha}\right)^{s_{1}}\times\frac{\rho_{1}(s_{1})^{s_{1}}}{1-\rho_{1}(s_{1})^{s_{1}}}+\frac{\rho_{1}(s_{1})}{1-\rho_{1}(s_{1})}\right),\Delta_{F_{0},T,F_{1}}\right\}.$ (S5)
@@
We only prove the case when $\tau$ is the CUSUM type stopping time $\tau_{\text{wcs-ripr}}$ defined as
$\tau_{\text{cs}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\prod_{i=j}^{n}\frac{p_{1}^{*}(X_{i})}{p_{0}^{*}(X_{i})}\geq A\right\}.$
Proofs for the other case, when $\tau$ is the SR-type stopping time, follow analogously.
$\mathbb{P}_{P,\infty}[\tau_{\text{cs}}\geq t]$ $=\mathbb{P}_{P,\infty}\left[\displaystyle\max_{1\leq k\leq t-1}\max_{1\leq j\leq k}\prod_{i=j}^{k}\frac{p_{1}^{*}(X_{i})}{p_{0}^{*}(X_{i})}<A\right].$
Since $(P_{0}^{*},P_{1}^{*})$ is an LFD pair for testing $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$, for any $P_{0}\in\mathcal{P}_{0},$ we have
${R(P_{0},\phi)\leq R(P_{0}^{*},\phi)}$
[…]
${\implies{P_{0}}\!\left(\frac{p_{1}^{*}(X)}{p_{0}^{*}(X)}>\eta\right)\leq{P_{0}^{*}}\!\left(\frac{p_{1}^{*}(X)}{p_{0}^{*}(X)}>\eta\right),\text{ for all }\eta}$
which is same as
$\frac{p_{1}^{*}(X)}{p_{0}^{*}(X)}\preceq_{\text{st}}\frac{p_{1}^{*}(X^{\prime})}{p_{0}^{*}(X^{\prime})},$
where $X\sim P$ and $X^{\prime}\sim P^{*}$. By Theorem 1.A.2 of [26], the stochastic ordering implies that there exist a random variable $Z$ and functions $\psi_{1}$ and $\psi_{2}$ such that $\psi_{1}(z)\leq\psi_{2}(z)$ for all $z$ and $\frac{p_{1}^{*}(X)}{p_{0}^{*}(X)}\stackrel{{\scriptstyle d}}{{=}}\psi_{1}(Z)$ and $\frac{p_{1}^{*}(X^{\prime})}{p_{0}^{*}(X^{\prime})}\stackrel{{\scriptstyle d}}{{=}}\psi_{2}(Z)$. Therefore,
$\mathbb{P}_{P,\infty}[\tau_{\text{cs}}\geq t]$ $=\mathbb{P}\left[\displaystyle\max_{1\leq k\leq t-1}\max_{1\leq j\leq k}\prod_{i=j}^{k}\psi_{1}(Z_{i})dW(\theta_{1})<A\right]$
$\geq\mathbb{P}\left[\displaystyle\max_{1\leq k\leq t-1}\max_{1\leq j\leq k}\prod_{i=j}^{k}\psi_{2}(Z_{i})dW(\theta_{1})<A\right]$
$=\mathbb{P}_{P^{*},\infty}[\tau_{\text{cs}}\geq t].$